OpenStudy (boobear7411):

(8x^4 + 4x^2 + x) - (x + 10)(2ax^2 + 1) The expression is simplified to 8x^4 - 6x^3 - 56x^2 - 10. What is the value of "a" ?

OpenStudy (amistre64):

you will have to expand it out and compare coefficients

OpenStudy (boobear7411):

and how do i do that ?

OpenStudy (amistre64):

the textbooks tend to use a method they call "foil"

OpenStudy (boobear7411):

whats that ?

OpenStudy (amistre64):

\[(a+b)(x+y)=ax+ay+bx+by\]

OpenStudy (amistre64):

(x + 10)(2ax^2 + 1) let a=x, b=10, x=2ax^2, and y = 1

OpenStudy (boobear7411):

Im so confused .

OpenStudy (amistre64):

its just multiplication ..... you can do it with some practice

OpenStudy (boobear7411):

okay

OpenStudy (amistre64):

(x + 10)(2ax^2 + 1) give this a shot, and ill see if i can correct any errors along the way

OpenStudy (boobear7411):

i have no idea whatsoever on how to do that . . .

OpenStudy (amistre64):

\[(a+b)(x+y)=ax+ay+bx+by\] let a=x, b=10, x=2ax^2, and y = 1 replace and multiply

OpenStudy (amistre64):

my choice of general variables is prolly bad seeing the it uses as and xs ... but try to see thru that little fauxpaux

OpenStudy (boobear7411):

(x+10)(2ax^2+1) . . . for the first part ? ? ?

OpenStudy (amistre64):

for the first part we would want to start expanding this with that yes \[(m+n)(p+q)=mp+mq+npx+nq\]let m=x, n=10, p=2ax^2, and q = 1

OpenStudy (boobear7411):

(x+10)(2ax^2+1) . . .

OpenStudy (amistre64):

im not going to be able to do it for you, and i really dont know what your asking when you restate: (x+10)(2ax^2+1)

OpenStudy (boobear7411):

can you at least show me the steps cause i have no idea how to do this . .

OpenStudy (amistre64):

i have shown you the steps. you multiply the first terms together, add that to the product of the first term in one, by the last term in the other, add that to the product of the last term in one by the first term in the other, and add that to the product of the last terms. which quite frankly is more confusing to write out than it is to read in notation:\[(m+n)(p+q)=mp+mq+np+nq\]

OpenStudy (amistre64):

ive got to head out to class, so ill chk back in on this later today to see if there is any progress.

OpenStudy (boobear7411):

yes .

OpenStudy (boobear7411):

i wouldnt know ?

OpenStudy (boobear7411):

?